An electric field with a strength of 1.25x104N/C exerted 2.50 N on a certain test charge, q. Determine the charge on q.

1
Expert's answer
2022-05-20T13:19:53-0400
"F=qE"

"q=\frac{F}{E}=\frac{2.50}{1.25*10^4}=2*10^{-4}\:\rm C"